OpenStudy (anonymous):

how do i solve : x^-5x+6 how do i solve : x^-5x+6 @Mathematics

OpenStudy (anonymous):

mult to be positive 6 6= 1*6 6=2*3 add to be -5 choose 1*6 (x-6)(x+1)

OpenStudy (anonymous):

THANK YOU !! :)

OpenStudy (anonymous):

do you mind showing me the steps please ?